# Graph is a Natural Regularization: Revisiting Vector Quantization for Graph Representation Learning

### v1

@@ Abstract
Vector Quantization (VQ) has recently emerged as a promising approach for learning discrete representations of graph-structured data. However, a fundamental challenge, i.e., codebook collapse, remains underexplored in the graph domain, significantly limiting the expressiveness and generalization of graph tokens. In this paper, we present the first empirical study showing that codebook collapse consistently occurs when applying VQ to graph data, even with mitigation strategies proposed in vision or language domains. To understand why graph VQ is particularly vulnerable to collapse, we provide a theoretical analysis and identify two key factors: early assignment imbalances caused by redundancy in graph features and structural patterns, and self-reinforcing optimization loops in deterministic VQ. To address these issues, we propose RGVQ, a novel framework that integrates graph topology and feature similarity as explicit regularization signals to enhance codebook utilization and promote token diversity. RGVQ introduces soft assignments via Gumbel-Softmax reparameterization, ensuring that all codewords receive gradient updates. In addition, RGVQ incorporates a structure-aware contrastive regularization to penalize the token co-assignments among similar node pairs. Extensive experiments demonstrate that RGVQ substantially improves codebook utilization and consistently boosts the performance of state-of-the-art graph VQ backbones across multiple downstream tasks, enabling more expressive and transferable graph token representations.
Introduction
@@
In recent years, a discretization-based tokenization method, known as Vector Quantization (VQ), has attracted significant research attention for its effectiveness in generative modeling (Van Den Oord, Vinyals et al. 2017; Caron et al. 2018). VQ quantizes continuous latent representations into discrete clusters referred to as “codewords” in a learnable codebook. These codewords are then trained to reconstruct the original data samples. By discretizing the latent space, VQ provides an effective prior for learning disentangled features, and has achieved remarkable success in various generative modeling tasks, including image synthesis (Ramesh et al. 2021; Chang et al. 2023), speech generation (Dhariwal et al. 2020), and language models (Liu et al. 2025; Van Baalen et al. 2024).
@@
Motivated by these successes, recent efforts have begun to explore the extension of VQ to graphs for scalable and versatile graph tokenization. First, discretizing graphs into VQ tokens enables compact graph compression, substantially reducing the memory and computation overhead during inference (Yang et al. 2023; Luo et al. 2024). Second, VQ provides a natural mechanism for abstracting structural patterns into a reusable token vocabulary, analogous to the language tokens used in Large Language Models (LLMs), and offers a promising pathway toward Graph Foundation Models (GFMs) (Wang et al. 2024). Third, VQ allows graphs to be serialized into token sequences, enabling sequence-based modeling with standard Transformer architectures that are widely adopted in NLP and vision, and eliminating the need for handcrafted inductive biases that are typically required in Graph Transformers (Wang et al. 2025).
@@
Observations. Similar to VQ models in the vision and language domains, which are typically trained to reconstruct input samples, Graph VQ is also trained with reconstruction objectives, i.e., node feature and link reconstruction. Nevertheless, through our empirical study, we observe that codebook collapse consistently occurs during the training, even when applying mitigation strategies that are commonly used in other domains. This refers to the phenomenon where most inputs are mapped to only a few codewords, leaving the majority underutilized (Zhu et al. 2024; Zhang et al. 2023). As a result, only a limited number of tokens can be utilized during inference, leading to overly coarse representations and significant degradation in performance. However, prior work focuses primarily on the performance of downstream tasks, without addressing this critical problem in the first place. This naturally leads to a central question: What are the root causes that make Graph VQ more prone to collapse?
@@
Challenges. To answer this question, we investigate the phenomenon from both empirical and theoretical perspectives. Empirically, we find that the severity of collapse is correlated with typical graph properties such as structural and feature redundancy, indicating that inherent characteristics of graph data can aggravate the issue. Based on this observation, we theoretically identify two underlying factors: (1) early assignment imbalances caused by similar features and local structures; and (2) self-reinforcing training dynamics in VQ models, where frequently assigned codewords receive more updates and become increasingly dominant, while rarely selected ones remain inactive. This amplifies assignment imbalances and ultimately leads to collapse.
@@
Solution. Based on these insights, we propose Regularized Graph VQ (RGVQ), a novel framework that integrates graph topology and feature similarity as explicit regularization signals to enhance codebook utilization. First, to break the self-reinforcing loops, RGVQ adopts the Gumbel-Softmax reparameterization to relax hard assignments into differentiable probability distributions, enabling the gradients to flow not only to the most likely codewords but also to the less probable candidates. Second, RGVQ leverages graph topology and feature similarity to regularize token assignment distributions, explicitly penalizing token co-assignments induced by graph redundancy. This regularization encourages nodes with similar features and local structures to share token distributions, while discouraging similar assignment among unrelated nodes. In summary, our main contributions can be briefed as follows.
@@
  • We perform comprehensive experiments on state-of-the-art (SOTA) Graph VQ backbones, demonstrating that our proposed method improves codebook utilization and serves as a general solution for graph token learning.
Related Work
Vector Quantization. Vector Quantization (VQ) maps continuous inputs to discrete tokens in a codebook and has been widely used in image, video, and audio generation (Chung, Tang, and Glass 2020). This success has motivated efforts to extend VQ to graph data. For example, VQ-GNN (Ding et al. 2021) and VQGraph (Yang et al. 2023) apply VQ in supervised node classification tasks for embedding compression. However, their fully supervised training deviates from the original unsupervised training scheme of VQ. More recently, GFT pretrains VQ by reconstructing graph features to utilize the learned codebook as transferable vocabulary across tasks and domains (Wang et al. 2024). GQT employs residual VQ to tokenize graphs for vanilla transformers, alleviating manual architectural bias in graph transformers (Wang et al. 2025). While both methods demonstrate promising applications of Graph VQ, they overlook the issue of codebook collapse, which undermines the generalization of learned tokens. HQA-GAE introduces a hierarchical VQ with annealing for codeword selection, improving performance on graph tasks (Zeng et al. 2025). Nevertheless, it does not effectively resolve the non-differentiability of VQ and lacks a formal analysis of codebook collapse.
Collapse Mitigation. One of the most fundamental limitations of VQ is codebook collapse, wherein only a small fraction of codewords are used. Various mitigation strategies have been explored in different domains. For example, Exponential Moving Average (EMA) is proposed to stabilize codebook updates. Pretraining the encoder (Zhao et al. 2024a) is proposed to mitigate embedding drift during training VQ. In addition, codebook reset (Zeghidour et al. 2021) periodically reinitializes inactive codewords with encoder embeddings. Affine parameters (Huh et al. 2023) introduce a learnable transformation before quantization to align encoder outputs with the codebook space and improve token utilization. Recently, SimVQ (Zhu et al. 2024) reparameterizes the code vectors through a linear transformation layer based on a learnable latent basis. Although these mitigation strategies have been evaluated in image and speech domains, their performance on graph data remains underexplored.
Preliminary
Graph Neural Network. Graph Neural Networks (GNNs) learn the node representations by recursively aggregating features from neighbors, also known as message-passing (Sun et al. 2022; Zhang et al. 2019). Formally, the representation of node $v$ at the $l$-th layer is:
@@
Deterministic VQ. VQ maps continuous representations into a finite set of discrete embeddings in the codebook (Van Den Oord, Vinyals et al. 2017). Given a codebook $\mathbf{C}=\{\mathbf{e}_{i}\}_{i=1}^{K}$ with each discrete codeword $\mathbf{e}_{i}\in\mathbb{R}^{d}$, a continuous input $h_{i}\in\mathbb{R}^{d}$ is quantized as $z_{i}$ with the nearest codebook vector $e_{k}$ by:
@@
where $\delta_{j}\in\{0,1\}^{1\times K}$ is the one-hot indicator vector with only the $j$-th element being 1. To enable gradient propagation through the non-differentiable vector $\delta_{j}$, the Straight-Through Estimator (STE) is applied (Bengio, Léonard, and Courville 2013). During the backward process, the gradient of the quantized embedding $z_{i}=\delta_{j}\mathbf{C}$ is copied to $h_{i}$, which is denoted as
@@
For Graph VQ, the reconstruction task typically involves reconstructing the graph properties, i.e., node features and links (Wang et al. 2024, 2025; Yang et al. 2023):
@@
where $A$ is the adjacency matrix, $N$ is the total number of nodes, $D(\cdot)$ is the decoder, and $X$ is the node feature matrix. Note that there are other forms of loss functions, e.g., utilizing the negative sampling to compute the link reconstruction loss (Wang et al. 2024; Zeng et al. 2025).
Metric for Codebook Utilization. The extent of codebook collapse is measured by the codebook perplexity (Takida et al. 2022; Yan et al. 2024), which is defined as:
@@
Empirical Study
We begin by investigating the codebook perplexity of Graph VQ on different graph datasets. Following the settings of prior work (Ding et al. 2021; Wang et al. 2024; Yang et al. 2023), we apply vanilla Graph VQ and its variants augmented with widely adopted collapse mitigation methods, including EMA, codebook reset, pretrained encoder, and affine parameters. By default, orthogonal normalization (Yu et al. 2021) and cosine similarities (Wang et al. 2024) are incorporated in all variants. More results and experiment settings can be found in Appendix A. From Figure 1, we make the following observations:
  • Codebook collapse is a systematic and severe issue in Graph VQ. Across all datasets, the perplexity of VQ remains far below the codebook capacity, and fails to grow proportionally with the increasing codebook size.
  • General mitigation strategies adopted from other domains only achieve marginal improvements and fail to fundamentally address codebook collapse in graphs.
These findings reveal that codebook collapse is not merely incidental, but a systematic issue in Graph VQ. We hypothesize that the unique properties of graph data, i.e., inherent feature redundancy and non-i.i.d. nature, may contribute to this phenomenon. To investigate this, we analyze two graph-level statistics that serve as proxies for these properties on different datasets. We consider: (1) PCA@95%, which quantifies feature redundancy by measuring the number of principal components needed to preserve 95% of node feature variance; and (2) average node degree, which reflects local connectivity density. Higher degrees imply stronger dependencies between neighboring nodes, violating the i.i.d. assumption and serving as a simple proxy for the non-i.i.d. nature of graph data (Yang et al. 2023). From Figure 2, we observe a positive correlation between PCA@95% and codebook perplexity, and a negative correlation between average degree and perplexity. Lower PCA@95% suggests higher feature redundancy, and higher average degree implies stronger local connectivity and non-i.i.d. characteristics, both associated with a greater tendency toward collapse.
Theoretical Analysis
@@
Given two nodes $v_{1}$ and $v_{2}$ and their computation tree $\mathcal{T}_{v_{1}}$ and $\mathcal{T}_{v_{2}}$, let $h_{v_{1}}=\phi(\mathcal{T}_{v_{1}})$, $h_{v_{2}}=\phi(\mathcal{T}_{v_{2}})$ be their embeddings under a GNN encoder $\phi$ with the distance between node embeddings $\Delta^{L}_{v_{1},v_{2}}=\|h_{v_{1}}-h_{v_{2}}\|_{2}$. The probability that two nodes are quantized to the same codeword $p:=\mathbb{P}[z_{v_{1}}=z_{v_{2}}]$ satisfies the following lower bound:
@@
The proof can be found in Appendix B. ∎
Remark 1.
Theorem 1 provides a theoretical interpretation of the empirical trends observed in Figure 2. When node features exhibit low variance or the graphs become more densely connected, the computation trees of different nodes become increasingly similar. This structural and feature redundancy increases the probability of token co-assignment, contributing to imbalanced codeword assignments.
Self-Reinforcing Dynamics of Graph VQ. Given the higher probability of token co-assignments induced by graph data, we further analyze how the optimization dynamics of Graph VQ progressively amplifies imbalances and ultimately leads to collapse. In Graph VQ, the codebook $\mathbf{C}$ is updated only through the vocabulary loss (Zhu et al. 2024), i.e., the second term in Equation 4. The update is denoted as:
@@
Gumbel-Softmax Reparameterization. In deterministic VQ, the training dynamics of hard assignments prevent gradient backpropagation to unselected codewords, ultimately leaving them inactive and underutilized. To address this issue, we adopt Gumbel-softmax reparameterization (Roy et al. 2018; Sønderby, Poole, and Mnih 2017), which replaces hard nearest-neighbor assignment with a differentiable soft selection. Formally, the assignment distribution of quantizing $h_{i}$ to entries in codebook $C$ is defined as:
@@
Structure-Aware Regularization. To mitigate token co-assignments induced by the redundancy of graph data, we incorporate feature and structural similarities to regularize the token assignment distribution, encouraging the model to avoid overuse of specific codebook entries. Our key insight is that while structurally or semantically related nodes should share similar representations, they should nonetheless be assigned to distinct tokens to preserve codebook diversity. Thus, we distinguish between similar and dissimilar node pairs based on feature and structural similarity. Formally, for a given anchor node $v$, we define:
  • Positive set $\mathcal{N}_{P}$: consists of the top-K nodes that are either structurally or semantically similar to $v$. Specifically, this includes nodes that satisfy at least one of the following conditions: (1) they are directly connected to $v$, i.e., $a_{uv}=1$, where $a_{uv}\in A$ is the adjacency matrix; or (2) they have high feature similarity to $v$, i.e., $\|x_{v}-x_{u}\|<\epsilon$, where $\epsilon$ is a predefined threshold.
  • Negative set $\mathcal{N}_{N}$: consists of the top-K nodes that are dissimilar to $v$ in both structure and feature. Specifically, these are nodes that simultaneously satisfy: (1) they are not connected to $v$ in the graph, i.e., $a_{uv}=0$; and (2) they exhibit low feature similarity to $v$, i.e., $\|x_{v}-x_{u}\|>\gamma$, where $\gamma>\epsilon$.
@@
We average $\mathcal{L}_{i}$ over all nodes to obtain the final regulation loss, i.e., $\mathcal{L}_{\text{reg}}=\sum_{N}\mathcal{L}_{i}$. This regularization term is then added to the loss function in Equation 4, forming the ultimate loss for training Graph VQ:

### v2

@@ Abstract
Vector Quantization (VQ) has recently emerged as a promising approach for learning discrete representations of graph-structured data. However, a fundamental challenge, i.e., codebook collapse, remains underexplored in the graph domain, significantly limiting the expressiveness and generalization of graph tokens. In this paper, we present the first empirical study showing that codebook collapse consistently occurs when applying VQ to graph data, even with mitigation strategies proposed in vision or language domains. To understand why graph VQ is particularly vulnerable to collapse, we provide a theoretical analysis and identify two key factors: early assignment imbalances caused by redundancy in graph features and structural patterns, and self-reinforcing optimization loops in deterministic VQ. To address these issues, we propose RGVQ, a novel framework that integrates graph topology and feature similarity as explicit regularization signals to enhance codebook utilization and promote token diversity. RGVQ introduces soft assignments via Gumbel-Softmax reparameterization, ensuring that all codewords receive gradient updates. In addition, RGVQ incorporates a structure-aware contrastive regularization to penalize the token co-assignments among dissimilar node pairs. Extensive experiments demonstrate that RGVQ substantially improves codebook utilization and consistently boosts the performance of state-of-the-art graph VQ backbones across multiple downstream tasks, enabling more expressive and transferable graph token representations.
1 Introduction
@@
In recent years, a discretization-based tokenization method, known as Vector Quantization (VQ), has attracted significant research attention for its effectiveness in generative modeling (Van Den Oord et al., 2017; Caron et al., 2018). VQ quantizes continuous latent representations into discrete clusters referred to as “codewords” in a learnable codebook (Zhang et al., 2023b). These codewords are then trained to reconstruct the original data samples. By discretizing the latent space, VQ provides an effective prior for learning disentangled features, and has achieved remarkable success in various generative modeling tasks, including image synthesis (Ramesh et al., 2021; Chang et al., 2023; Li et al., 2024a), speech generation (Dhariwal et al., 2020; Zhang et al., 2024), and language models (Liu et al., 2025; Van Baalen et al., 2024).
@@
Motivated by these successes, recent efforts have begun to explore the extension of VQ to graphs for scalable and versatile graph tokenization. First, discretizing graphs into VQ tokens enables compact graph compression, substantially reducing the memory and computation overhead during inference (Yang et al., 2023; Luo et al., 2024). Second, VQ provides a natural mechanism for abstracting structural patterns into a reusable token vocabulary, analogous to the language tokens used in Large Language Models (LLMs), and offers a promising pathway toward Graph Foundation Models (GFMs) (Wang et al., 2024). Third, VQ allows graphs to be serialized into token sequences, enabling sequence-based modeling with standard Transformer architectures that are widely adopted in NLP and vision, and eliminating the need for handcrafted inductive biases that are typically required in Graph Transformers (Wang et al., 2025).
@@
Similar to VQ models in the vision and language domains, which are typically trained to reconstruct input samples (Navaneet et al., 2024; Deng et al., 2024), Graph VQ is also trained with reconstruction objectives, i.e., node feature and link reconstruction. Nevertheless, through our empirical study, we observe that codebook collapse consistently occurs during the training, even when applying mitigation strategies that are commonly used in other domains. This refers to the phenomenon where most inputs are mapped to only a few codewords, leaving the majority underutilized (Zhu et al., 2024; Zhang et al., 2023a). As a result, only a limited number of tokens can be utilized during inference, leading to overly coarse representations and significant degradation in performance. However, prior work focuses primarily on the performance of downstream tasks, without addressing this critical problem in the first place (Wang et al., 2024; Zeng et al., 2025; Yang et al., 2023). This consistent underutilization naturally raises a central question: What makes Graph VQ more prone to collapse?
@@
To answer this question, we investigate the phenomenon from both empirical and theoretical perspectives. Empirically, we find that the severity of collapse is correlated with typical graph properties such as structural and feature redundancy, indicating that inherent characteristics of graph data can aggravate the issue. Based on this observation, we theoretically identify two underlying factors: (1) early assignment imbalances caused by similar features and local structures; and (2) self-reinforcing training dynamics in VQ models, where frequently assigned codewords receive more updates and become increasingly dominant, while rarely selected ones remain inactive. This amplifies assignment imbalances and ultimately leads to collapse.
@@
Based on these insights, we propose Regularized Graph Vector Quantization (RGVQ), a novel framework that integrates graph topology and feature similarity as explicit regularization signals to enhance codebook utilization. First, to break the self-reinforcing loops, RGVQ adopts the Gumbel-Softmax reparameterization to relax hard assignments into differentiable probability distributions, enabling the gradients to flow not only to the most likely codewords but also to the less probable candidates. Second, RGVQ leverages graph topology and feature similarity to regularize token assignment distributions, explicitly penalizing token co-assignments induced by graph redundancy. This regularization encourages nodes with similar features and local structures to share token distributions, while discouraging similar assignment among unrelated nodes. Our contributions can be summarized as follows.
@@
  • We perform comprehensive experiments on state-of-the-art (SOTA) Graph VQ backbones, demonstrating that our proposed method improves codebook utilization and downstream performance, and serves as a general solution for graph token learning.
2 Related Work
Vector Quantization. Vector Quantization (VQ) maps continuous inputs to discrete tokens in a codebook and has been widely used in image, video, and audio generation (Chung et al., 2020; Fifty et al., 2024; Tang et al., 2022). This success has motivated efforts to extend VQ to graph data. For example, VQ-GNN (Ding et al., 2021) and VQGraph (Yang et al., 2023) apply VQ in supervised node classification tasks for embedding compression. However, their fully supervised training deviates from the original unsupervised training scheme of VQ (Chen & Lee, 2021; Yu et al., 2021b). More recently, GFT pretrains VQ by reconstructing graph features to utilize the learned codebook as transferable vocabulary across tasks and domains (Wang et al., 2024). GQT employs residual VQ to tokenize graphs for vanilla transformers, alleviating manual architectural bias in graph transformers (Wang et al., 2025). While both methods demonstrate promising applications of Graph VQ, they overlook the issue of codebook collapse, which undermines the generalization of learned tokens. HQA-GAE introduces a hierarchical VQ with annealing for codeword selection, improving performance on graph tasks (Zeng et al., 2025). Nevertheless, it does not effectively resolve the non-differentiability of VQ and lacks a formal analysis of codebook collapse.
Collapse Mitigation. One of the most fundamental limitations of VQ is codebook collapse, wherein only a small fraction of codewords are used (Zhang et al., 2024; Lu et al., 2023). Various mitigation strategies have been explored. For example, Exponential Moving Average (EMA) is proposed to stabilize codebook updates (Polyak & Juditsky, 1992; Wu & Yu, 2019). Pretraining the encoder (Zhao et al., 2024) is proposed to mitigate embedding drift during training VQ. In addition, codebook reset (Zeghidour et al., 2021; Williams et al., 2020) periodically reinitializes inactive codewords with encoder embeddings. Affine parameters (Huh et al., 2023; Zhang et al., 1997) introduce a learnable transformation before quantization to align encoder outputs with the codebook space. Recently, SimVQ (Zhu et al., 2024) reparameterizes the code vectors through a linear transformation layer based on a learnable latent basis. Although these mitigation strategies have been evaluated in image and speech domains, their performance on graph data remains underexplored.
3 Preliminary
Graph Neural Network. Graph Neural Networks (GNNs) learn the node representations by recursively aggregating features from neighbors, also known as message-passing (Sun et al., 2022; Zhang et al., 2019). Formally, the representation of node $v$ at the $l$-th layer is:
@@ 3 Preliminary
Deterministic VQ. VQ maps continuous vectors into a finite set of discrete embeddings in the codebook (Van Den Oord et al., 2017). Given a codebook $\mathbf{C}=\{\mathbf{e}_{i}\}_{i=1}^{K}$ with each discrete codeword $\mathbf{e}_{i}\in\mathbb{R}^{d}$, a continuous input $h_{i}\in\mathbb{R}^{d}$ is quantized as $z_{i}$ with the nearest codebook vector $e_{k}$ by:
@@
where $\delta_{j}\in\{0,1\}^{1\times K}$ is the one-hot indicator vector with only the $j$-th element being 1. To enable gradient propagation through the non-differentiable vector $\delta_{j}$, the Straight-Through Estimator (STE) is applied (Bengio et al., 2013). During the backward process, the gradient of the quantized embedding $z_{i}=\delta_{j}\mathbf{C}$ is copied to $h_{i}$, which is denoted as
@@
For Graph VQ, the reconstruction task typically involves reconstructing the graph properties, i.e., node features and links (Wang et al., 2024; 2025; Yang et al., 2023):
@@
where $A$ is the adjacency matrix, $N$ is the total number of nodes, $D(\cdot)$ is the decoder, and $X$ is the node feature matrix. Note that there are other forms of loss functions, e.g., utilizing the negative sampling to compute the link reconstruction loss (Wang et al., 2024; Zeng et al., 2025).
Metric for Codebook Utilization. The extent of codebook collapse is measured by the codebook perplexity (Takida et al., 2022; Yan et al., 2024; Zheng & Vedaldi, 2023), which is defined as:
@@ 4 Motivation
4.1 Empirical Study
We begin by investigating the codebook perplexity of Graph VQ on different graph datasets. Following the settings of prior work (Ding et al., 2021; Wang et al., 2024; Yang et al., 2023), we apply vanilla Graph VQ and its variants augmented with widely adopted collapse mitigation methods, including EMA, codebook reset, pretrained encoder, and affine parameters. By default, orthogonal normalization (Yu et al., 2021a) and cosine similarities (Wang et al., 2024) are incorporated in all variants. The implementation details and additional datasets can be found in Appendix B and C, respectively. From Figure 1, we make the following observations: (Ob. 1) Codebook collapse is a systematic and severe issue in Graph VQ. Across all datasets, the perplexity of VQ remains far below the codebook capacity, and fails to grow proportionally with the increasing codebook size. (Ob. 2) General mitigation strategies adopted from other domains only achieve marginal improvements and fail to fundamentally address codebook collapse in graphs.
These findings reveal that codebook collapse is not merely incidental, but a systematic issue in Graph VQ. We hypothesize that the unique properties of graph data, i.e., inherent feature redundancy and non-i.i.d. nature, may contribute to this phenomenon. To investigate this, we analyze two graph-level statistics that serve as proxies for these properties on investigated datasets. We consider: (1) PCA@95%, which quantifies feature redundancy by measuring the number of principal components needed to preserve 95% of node feature variance (Dong et al., 2022; Hou et al., 2023); and (2) average node degree, which reflects local connectivity density. Higher degrees imply stronger dependencies between neighboring nodes, violating the i.i.d. assumption and serving as a simple proxy for the non-i.i.d. nature of graph data (Yang et al., 2023; Wu et al., 2019). From Figure 2, we observe a positive correlation between PCA@95% and codebook perplexity, and a negative correlation between average degree and perplexity. Lower PCA@95% suggests higher feature redundancy, and higher average degree implies stronger local connectivity and non-i.i.d. characteristics, both associated with a greater tendency toward collapse.
4.2 Theoretical Analysis
@@
Given two nodes $v_{1}$ and $v_{2}$ and their corresponding computation trees $\mathcal{T}_{v_{1}}$ and $\mathcal{T}_{v_{2}}$ under a message-passing GNN encoder $\phi$, let $h_{v_{1}}=\phi(\mathcal{T}_{v_{1}})$, $h_{v_{2}}=\phi(\mathcal{T}_{v_{2}})$ be their embeddings with the distance between node embeddings $\Delta^{L}_{v_{1},v_{2}}=\|h_{v_{1}}-h_{v_{2}}\|_{2}$. The probability that two nodes are quantized to the same codeword $p:=\mathbb{P}[z_{v_{1}}=z_{v_{2}}]$ satisfies the following lower bound:
$p\geq\Bigg(1-\frac{1}{\delta_{c}}\Big[C_{1}\|x_{v_{1}}-x_{v_{2}}\|_{2}+\ C_{2}\sum_{j\in\mathcal{N}(v)}\Delta^{L-1}_{v_{1},v_{2},j}\Big]\Bigg)\geq\Bigg(1-\frac{2B_{x}}{\delta_{c}}\Big[C_{1}+\sum_{\ell=1}^{L}C_{2}^{\ell}D_{\ell}\Big]\Bigg),$ (7)
@@
The proof can be found in Appendix A. ∎
Remark 1. Theorem 1 provides a theoretical interpretation of the empirical trends observed in Figure 2. When node features exhibit low variance or the graphs become more densely connected, the computation trees of different nodes become increasingly similar. This structural and feature redundancy increases the probability of token co-assignment, contributing to imbalanced codeword assignments.
Self-Reinforcing Dynamics of Graph VQ. Given the higher probability of token co-assignments induced by graph data, we further analyze how the optimization dynamics of Graph VQ progressively amplifies imbalances and ultimately leads to collapse. In Graph VQ, the codebook $\mathbf{C}$ is updated only through the vocabulary loss (Zhu et al., 2024), i.e., the second term in Equation 4. The update is denoted as:
@@ 5 Methodology
Gumbel-Softmax Reparameterization. In deterministic VQ, the training dynamics of hard assignments prevent gradient backpropagation to unselected codewords, ultimately leaving them inactive and underutilized. To address this issue, we adopt Gumbel-softmax reparameterization (Roy et al., 2018; Sønderby et al., 2017), which replaces hard nearest-neighbor assignment with a differentiable soft selection. Formally, the assignment distribution of quantizing $h_{i}$ to entries in codebook $C$ is defined as:
@@
Structure-Aware Regularization. To mitigate token co-assignments induced by the redundancy of graph data, we incorporate feature and structural similarities to regularize the token assignment distribution, encouraging the model to avoid overuse of specific codebook entries. Our key insight is that collapse arises when nodes are spuriously mapped to the same tokens due to overly similar computation-tree embeddings caused by structural or feature redundancy. Therefore, we explicitly distinguish between similar and dissimilar node pairs based on both feature and structural similarity: similar nodes can exhibit more consistent assignment distributions, while dissimilar nodes should be discouraged from co-assignments. Formally, for a given anchor node $v$, we define:
  • Positive set $\mathcal{N}_{P}$: consists of the $n$ sampled nodes that are either structurally or semantically similar to $v$. Specifically, $n$ positive nodes are sampled from the union of the following two candidate sets: (1) nodes directly connected to $v$; or (2) the top $K$ feature-similar nodes to $v$. Formally, the positive set is denoted as:
    $\mathcal{N}_{P}=\left\{\,u\;\middle|\;\big(a_{uv}=1\big)\;\lor\;\Big(u\in\operatorname{arg\,topk}_{u^{\prime}\in\mathcal{V}}\text{sim}(x_{u},x_{v})\Big);\forall v\in\mathcal{V}\,\right\},$ (13)
    where $a_{uv}\in A$ is the adjacency matrix, sim(·,·) is the similarity function, $x_{u}$ is the feature vector of node $u$. We apply the cosine similarity as the similarity function.
  • Negative set $\mathcal{N}_{N}$: consists of the $n$ sampled nodes that are neither structurally connected nor semantically similar to $v$. Formally, the negative set is defined as:
    $\mathcal{N}_{N}=\left\{\,u\;\middle|\;(a_{uv}=0)\;\land\;\Big(u\notin\operatorname{arg\,topk}_{u^{\prime}\in\mathcal{V}}\text{sim}(x_{u},x_{v})\Big)\;\land\;(u\neq v),\;\forall v\in\mathcal{V}\,\right\}.$ (14)
@@
We sum $\mathcal{L}_{i}$ all nodes to obtain the final regulation loss, i.e., $\mathcal{L}_{\text{reg}}=\sum_{N}\mathcal{L}_{i}$. This regularization term is then added to the reconstruction loss in Equation 4, forming the ultimate loss:
@@
6 Experiments
To evaluate the generality and effectiveness of RGVQ, we conduct extensive experiments based on three key functions of Graph VQ: (1) codebook utilization, (2) transferability, and (3) serialization. First, we evaluate the codebook utilization of RGVQ on both homophilous and heterophilous graphs, aiming to verify its ability to mitigate codebook collapse. Second, we investigate the transferability. We integrate RGVQ into GFT (Wang et al., 2024), a graph foundation model that utilizes the learned codebook as pretrained graph tokens, and evaluate the performance on cross-task and cross-domain graphs. Third, we assess the serialization capability of RGVQ by examining its compatibility with sequence-based models. We use GQT (Wang et al., 2025), a transformer taking VQ tokens as input sequences, and evaluate the performance on node classification. Detailed dataset statistics, baselines, implementation details, and the complexity analysis are provided in Appendix B and D, respectively.